# Towards Inducing Document-Level Abilities in Standard Multilingual Neural Machine Translation Models

### v2

@@ Abstract
Neural Machine Translation (NMT) models have traditionally used Sinusoidal Positional Embeddings (PEs), which often struggle to capture long-range dependencies and are less efficient for handling extended context or document-level translation tasks. This work addresses the challenge of transitioning pre-trained NMT models from absolute sinusoidal PEs to relative PEs, such as Rotary Positional Embeddings (ROPE) and Attention with Linear Biases (ALIBI), without compromising performance. We demonstrate that parameter-efficient fine-tuning, using only a small amount of high-quality data, can successfully facilitate this transition. Experimental results indicate that switching from sinusoidal to relative PEs results in competitive translation quality on sentence-level evaluation benchmarks. Additionally, models trained with ROPE consistently outperform those using ALIBI and Sinusoidal PEs on document-level benchmarks across both string-based metrics and qualitative evaluations. Moreover, we find that a small amount of long-context data in a few languages is sufficient for cross-lingual length generalization, thereby inducing long-context capabilities.
Towards Inducing Document-Level Abilities in Standard Multilingual Neural Machine Translation Models
@@ 1 Introduction
In this work, we investigate whether it is possible to replace the positional embeddings in pre-trained NMT models with relative positional embeddings post-hoc. Our study explores both parameter-efficient fine-tuning (PEFT) methods and full fine-tuning (FFT) strategies to restore model performance while enabling long-context generalization.
In the remainder of this paper, we present the following contributions:
@@
  • Lastly, we open-source111 https://github.com/VarunGumma/fairseq our framework to support the efficient training and fine-tuning of long-context machine translation models.
@@ 2 Related Works
Zhang et al. (2019); Firat et al. (2016); Aharoni et al. (2019); Arivazhagan et al. (2019) has emerged as a leading methodology for creating translation systems capable of managing multiple languages. These MNMT systems leverage a unified encoder-decoder transformer framework Vaswani et al. (2017) along with language-specific embeddings to integrate linguistic nuances. Previous works Fan et al. (2021); Costa-jussà et al. (2022) have developed many-to-many (X-Y) models at a global scale, while other works Ramesh et al. (2022); Gala et al. (2023) have primarily focused on many-to-one (X-En) and one-to-many (En-X) multilingual models for Indic languages. An in-depth examination of MNMT methodologies is presented in the survey by Dabre et al. (2020).
@@
Transformer-based models require an explicit Positional Embedding (PE) information as all the timesteps are processed parallely, and hence there is a need for explicit positional. To this end, Vaswani et al. (2017) propose an empirical absolute PE approach following a fourier series with alternating sine and cosine terms. Vaswani et al. (2017) also experiment with Learnable PEs, however they do not show a significant performance difference as compared to absolute Sinusoidal Embeddings. Later works Rosendahl et al. (2019); Shaw et al. (2018) have demonstrated the advantages of relative PEs in transformer. Press et al. (2022) discuss ALIBI, a simple relative PE strategy achieved by adding the relative distance between tokens to the attention matrix and show that it can help the model generalize to longer sequence during inference even though trained on shorter sequences. Rotary Positional Embeddings (ROPE) Su et al. (2024) introduces the concept of rotating vectors according to their positions to explicitly encode relative position dependency in the self-attention formulation. Recent works Sun et al. (2023); Peng et al. (2023); Ding et al. (2024); Ostmeier et al. (2024) have also provided fixes to adapt ROPE to much longer sequences. In this work, we restrict ourselves to the standard variant of ROPE however, our method is generalizable to any other variant as well. In this study, we focus on the standard variant of RoPE, but our approach would be generalizable to other variants as well.
@@
Houlsby et al. (2019); Bapna and Firat (2019); Hu et al. (2022) deals with the concept of updating only a certain subset of parameters of the network, to efficiently generalize the model to new tasks or domains, instead of tuning all parameters, which might lead to catastrophic forgetting or over-fitting in large models. Low-Rank Adaptation (LORA) Hu et al. (2022) is currently the most popular parameter efficient finetuning strategy, which hypothesizes that the update to weights are low-rank, and injects small trainable parameters directly into the pre-trained weights to mimic the low-rank decomposed updates.
@@
Evaluation in machine translation is a well-explored area, with string-based metrics like BLEU (Papineni et al., 2002), ChrF (Popović, 2015), and ChrF++ (Popović, 2017) being the most commonly used. However, prior studies (Kocmi et al., 2021) have pointed out the limitations of BLEU, noting its bias towards exact matches, which makes it less suitable for languages with flexible word order. Trained metrics, such as COMET (Rei et al., 2020, 2022) and BLEURT (Sellam et al., 2020), utilize a regressor on top of an encoder to predict translation quality. These metrics have been shown to correlate more closely with human judgments, outperforming traditional string-based measures (Kocmi et al., 2021). However, their performance can vary significantly across different languages, and may not always be reliable for Indian languages (Gala et al., 2023).
@@
Lately, LLMs are emerging as general-purpose evaluators in Multilingual scenarios Hada et al. (2024b, a), exhibiting superior correlations with human judgment. Similar approaches have also been applied to the MT task as well. GEMBA (Kocmi and Federmann, 2023b) was one of the initial attempts to use LLMs to evaluate machine translation outputs. Follow up works like EAPROMPT (Lu et al., 2024), AutoMQM (Fernandes et al., 2023), GEMBA-MQM (Kocmi and Federmann, 2023a) have used LLMs to evaluate translations in a Multidimensional Quality Metrics (MQM) framework (Lommel et al., 2014; Freitag et al., 2021), by prompting them to identiy and penalize erroneous spans of text offering more interpretable and standardized evaluation. These appraoches demonstrate superior correlation with human judgements when compared to other model-based metrics like COMET, BLEURT. Given that LLMs typically support longer context lengths, they are well-suited for evaluating tasks involving extended input sequences. Consequently, inspired by GEMBA-MQM (Kocmi and Federmann, 2023a) we employ GPT-4o as an evaluator to investigate the qualitative differences between these baseline models.
@@ 3.1 Model
Our experimentation focuses on standard Encoder-Decoder transformer models. While several massively multilingual models, such as NLLB (Costa-jussà et al., 2022) and its distilled variants, have been released, our experimentation requires the availability of long-context or document-level training and evaluation datasets. Consequently, we focus on language-family specific models that allow for more controlled experimentation. For our experiments, we chose the distilled variants of IndicTrans2 Gala et al. (2023); Gumma et al. (2023). A key difference is that the distilled models released in Gala et al. (2023) we used were not trained with BackTranslation Sennrich et al. (2016) data, which we incorporate in our experimental setup. We experimented with models based on ROPE amd ALIBI by replacing the sinusoidal Positional Embedding in the pre-trained model with the corresponding relative Positional Embedding. This replacement led to a significant performance drop, prompting us to fine-tune the models using high-quality data Mohiuddin et al. (2022); Gumma et al. (2023); Gala et al. (2023) to study their recovery in each setting.
@@ 3.2.1 Train set
For maximum performance gains for our models, we choose a high-quality mixture of seen and unseen distributions for our training data. To this end, for the seen distribution, we select the BPCC-Seed, which is a high-quality human-verified subset of BPCC (Bharat Parallel Corpus Collection), and has shown to be effective in improving model performance Gala et al. (2023). As for the unseen distribution, we sample the top 150K pairs for each language from the BackTranslation (BPCC-BT) subset of BPCC based on the provided LASER Costa-jussà et al. (2022) or LaBSE Feng et al. (2022) cosine-similarity scores222whichever was available for that language. In case similarity scores are unavailable for a translation pair, a random sample was chosen. Additionally, for the unseen data, we also include the Asian Language Treebank [^3] https://www2.nict.go.jp/astrec-att/member/mutiyama/ALT (ALT) Corpus Riza et al. (2016), and the CoPara Aligned corpus E et al. (2023), which contains document-level translation pairs for 4 Dravidian languages, Telugu, Kannada, Malayalam and Tamil. The overall statistics of the dataset is available in table 1.
@@
Table 2: Distribution of long-context data in the fine-tuning data mixture. Indic-En pairs also follow the same distribution.
@@ 3.2.2 Dev/Test sets
Most existing benchmarks for Indian languages are composed of sentence-level test sets, which may not adequately reflect long-context performance or extrapolation capabilities. To address this, we construct document-level development and test sets using Flores [^4] https://huggingface.co/datasets/facebook/Flores (Costa-jussà et al., 2022) and IN22-Conv [^5] https://huggingface.co/datasets/ai4bharat/IN22-Conv (Gala et al., 2023). In this setup, sentences are grouped within the same context window, making them suitable for document-level evaluation. For Flores, we group sentences by metadata, including URL, domain, and topic headers, to map sentence-level pairs into document-level pairs, where each document-level unit consists of three sentences. This method is viable since this benchmark was created by professional translators who translate three consecutive sentences from every selected context window, as described by Goyal et al. (2022). As a result, we obtain 286 merged sentence pairs per language, with a total of 5,720 pairs. In case of IN22-Conv, the dataset was originally created by translating entire conversations from English into the target languages by professional translators, and used for sentence-level evaluation by Gala et al. (2023). We use metadata provided to trace back the conversations and treat the entire conversation as a sequence for long-context evaluation. This dataset has 44 conversations per language and covers 22 languages. The token length distribution of the synthesized IN22-Conv test set is shown in Figure 1. IN22-Conv serves as an accurate representation of a true document-level evaluation set due to its contextual dependencies across multiple turns in the conversation. Additionally, it supports all the languages under consideration, making it the primary benchmark for document-level evaluation.
@@ 3.3 Metrics and Evaluation
For consistency, we use the ChrF++ score (Popović, 2017) as the primary string-based metric, and follow the same evaluation practices as outlined in Gala et al. (2023) by normalizing the indic texts using indic_nlp_library [^6] https://github.com/anoopkunchukuttan/indic_nlp_library. For brevity, we report the aggregate score across all the languages in the respective test set, which are denoted in Table 3. For GPT-4o-based evaluation, we use the prompt template from GEMBA-MQM (Kocmi and Federmann, 2023a) as a starting point and make slight adjustments to prompt the model to additionally provide an implicit quality score between 0 and 1, in addition to identifying the error spans follwing MQM guidelines. Another notable difference is that, since we are evaluating long-context translations, including few-shot examples would be inefficient due to the limited context length. Therefore, we opt for a zero-shot evaluation setup. We report both the implicit score as rated by GPT-4o as well as calculated MQM-score (based on the scoring guidelines outlined in Kocmi and Federmann (2023a). We start with an ideal score of 25, deducting points based on identified errors and their severity: critical errors are penalized by a factor of 5, major by 3, and minor by 1. The final score is normalized by dividing by 25, yielding a value between 0 and 1.
@@ 3.4 Finetuning Strategies
First, we choose a full-fine-tuning (FFT) approach which tweaks all the 211.78M parameters of the model. This would allow the embeddings, layer-normalizations, and the rest of the linear layers777self-attention, cross-attention, feed-forward modules to be tuned as per the new relative Positional Embedding.
@@ 4 Experiments
We use the fairseq [^8] Our variant implements, LORA, ROPE, ALIBI, and a slightly efficient SDPA as per the requirements of this work framework Ott et al. (2019) for all of our experiments. All the finetuning experiments were conducted on 8 $\times$ H100 80Gb GPUs, and every model was trained till convergence with early-stopping. The exact hyperparameter sets are mentioned in Table 4.
@@
Patience 10
Max Positions (4096, 4096)
Warmup steps 2000
@@ 5 Results and Discussions
Figure 2: ChrF++ scores on the Document-level benchmarks of baselines involving 3 fine-tuning setups FFT, LORA, min-LORA comparing 3 types of postional embeddings namely SINE, ALIBI, ROPE along with the pre-trained IT2 performance baseline.
Figure 3: ChrF++ scores on the Sentence-level benchmarks of baselines involving 3 fine-tuning setups FFT, LORA, min-LORA comparing 3 types of postional embeddings namely SINE, ALIBI, ROPE along with the pre-trained IT2 performance baseline.
@@ 5.1 Swapping out Positional Embeddings
Figures 3 and 2 demonstrate the performance of several experimental setups that involve swapping of Positional Embeddings and followed by fine-tuning (if applicable). Detailed results are also presented in Tables 5 and 6.
@@
A notable finding in this setting is that fine-tuning a model originally using sinusoidal embeddings yielded performance comparable to models fine-tuned after the switch to using ROPE or ALIBI embeddings on sentence-level benchmarks as observed in Figure 3. This further reinforces the conclusion that Positional Embeddings in pre-trained transformer models can be safely tweaked even in later stages. Although sinusoidal, rotary, and ALIBI embeddings show similar performance in sentence-level test sets. The advantage of switching to ROPE or ALIBI can be observed in case of document-level test sets in line with the findings of Su et al. (2024) and Press et al. (2022).
@@
Our fine-tuning dataset primarily comprised sentence-level data, as detailed in Section 3.2.1, with only a limited amount of long-context or paragraph-level data available for four Dravidian languages accounting for less than 0.05% of the complete fine-tuning data mixture (see Table 1). The token distribution across this dataset is illustrated in Table 2. While most languages lacked long-context data, evaluation on document-level benchmarks still demonstrated strong performance on these, as seen in Figure 2. This indicates that fine-tuning with a small amount of long-context data introducing relative positional embeddings can enable robust cross-lingual length generalization. This finding is particularly important because long-context data is often scarce across many languages. Nevertheless, effective adaptation to longer contexts was achieved through cross-lingual generalization, even with minimal long-context data in select languages.
@@
ROPE has lower throughput compared to Sinusoidal embeddings due to its higher number of operations. Since the rotation function in ROPE is applied at every layer, it has a time complexity of $\mathcal{O}(n)$, whereas Sinusoidal embeddings are added only once, after the token embeddings, resulting in $\mathcal{O}(1)$ complexity. The exact decoding rates (in tokens per second) are illustrated in Figure Figure 5. An efficient implementation of ALIBI Press et al. (2022) achieves $\mathcal{O}(1)$ complexity, making it slightly slower than Sinusoidal but still significantly faster than ROPE. Therefore, for a basic sentence-level MT system, Sinusoidal PEs offer the best efficiency in terms of throughput and performance. However, for document-level or longer contexts, ROPE provides notable qualitative advantages, although it comes with a decrease in throughput.

### v3

# Towards Inducing Long-Context Abilities in Multilingual Neural Machine Translation Models

@@ Abstract
Neural Machine Translation (NMT) models have traditionally used Sinusoidal Positional Embeddings (PEs), which often struggle to capture long-range dependencies and are inefficient for handling extended context or document-level translation tasks. This work addresses the challenge of transitioning pre-trained NMT models from absolute Sinusoidal PEs to Relative PEs, such as RoPE and ALiBi, without compromising performance. We demonstrate that parameter-efficient fine-tuning, using only a small amount of high-quality data, can successfully facilitate this transition. Experimental results indicate that switching from Sinusoidal to Relative PEs results in competitive translation quality on sentence-level evaluation benchmarks. Additionally, models trained with RoPE consistently outperform those using ALiBi and Sinusoidal PEs on document-level benchmarks across both string-based metrics and qualitative evaluations. Moreover, we find that a small amount of long-context data in a few languages is sufficient for cross-lingual length generalization, thereby inducing long-context capabilities.
Towards Inducing Long-Context Abilities in
Multilingual Neural Machine Translation Models
@@ 1 Introduction
In this work, we investigate whether it is possible to replace the positional embeddings in pre-trained NMT models with relative positional embeddings post-hoc. Our study explores both parameter-efficient fine-tuning (PEFT) methods and full fine-tuning (FFT) strategies to restore model performance while enabling long-context generalization. In the remainder of this paper, we present the following contributions:
@@
  • Lastly, we open-source111 https://github.com/VarunGumma/fairseq our framework to support the efficient training and fine-tuning of long-context machine translation models. Further, our best-performing RoPE models are also available on HuggingFace222 https://aka.ms/IndicTrans2-RoPE .
@@ 2 Related Works
Zhang et al. (2019); Firat et al. (2016); Aharoni et al. (2019); Arivazhagan et al. (2019) has emerged as a leading methodology for creating translation systems capable of managing multiple languages. These MNMT systems leverage a unified encoder-decoder transformer framework Vaswani et al. (2017), along with language-specific embeddings to incorporate linguistic nuances. Previous works Fan et al. (2021); Costa-jussà et al. (2022) have developed many-to-many models at a global scale, while other works Ramesh et al. (2022); Gala et al. (2023) have primarily focused on many-to-one and one-to-many multilingual models for Indic languages. An in-depth examination of MNMT methodologies is presented in the survey by Dabre et al. (2020).
@@
Transformer-based models require explicit positional information, as all timesteps are processed in parallel. To address this, Vaswani et al. (2017) proposed an empirical absolute PE approach, using a Fourier series with alternating sine and cosine terms. Vaswani et al. (2017) also experimented with learnable PEs, but these did not show significant performance differences compared to absolute Sinusoidal embeddings. Later works Rosendahl et al. (2019); Shaw et al. (2018) demonstrated the advantages of relative PEs in transformers. Press et al. (2022) discussed ALiBi, a simple Relative PE strategy that adds relative token distances to the attention matrix, enabling models to generalize to longer sequences during inference, even when trained on shorter sequences. Rotary Positional Embeddings (RoPE) Su et al. (2024) introduced the concept of rotating vectors according to their positions to encode relative positional dependencies directly into the self-attention mechanism. Recent works Sun et al. (2023); Chen et al. (2023); Peng et al. (2023); Ding et al. (2024); Ostmeier et al. (2024) have also proposed adaptations of RoPE for much longer sequences. In this study, we focus on the standard variant of RoPE, though our method is generalizable to other variants.
@@
Houlsby et al. (2019); Bapna and Firat (2019); Hu et al. (2022) focuses on updating only a subset of model parameters to efficiently generalize to new tasks or domains, rather than fine-tuning all parameters. This approach mitigates the risks of catastrophic forgetting or overfitting in large models. Low-Rank Adaptation (LoRA) Hu et al. (2022), the most popular parameter-efficient fine-tuning strategy, hypothesizes that weight updates are low-rank and injects small trainable parameters into pre-trained weights to approximate low-rank updates.
@@
Evaluation in machine translation is well explored, with string-based metrics like BLEU (Papineni et al., 2002), ChrF (Popović, 2015), and ChrF++ (Popović, 2017) being the most commonly used. However, prior studies (Kocmi et al., 2021) have pointed out limitations of BLEU, particularly its bias towards exact matches, which makes it less suitable for languages with flexible word order. Trained metrics, such as COMET (Rei et al., 2020, 2022) and BLEURT (Sellam et al., 2020), employ a regressor on top of an encoder to predict translation quality. These metrics have been shown to correlate more closely with human judgments, outperforming traditional string-based measures (Kocmi et al., 2021). However, their performance can vary significantly across different languages and may not always be reliable for Indian languages (Gala et al., 2023).
@@
Recently, large language models (LLMs) have emerged as general-purpose evaluators in multilingual scenarios Hada et al. (2024b, a), exhibiting superior correlation with human judgment. This approach has also been applied to machine translation (MT). GEMBA (Kocmi and Federmann, 2023b) was one of the initial attempts to use LLMs for evaluating MT outputs. Follow-up works like EAPROMPT (Lu et al., 2024), AutoMQM (Fernandes et al., 2023), and GEMBA-MQM (Kocmi and Federmann, 2023a) have utilized LLMs in a Multidimensional Quality Metrics (MQM) framework (Lommel et al., 2014; Freitag et al., 2021), prompting models to identify and penalize erroneous spans of text, thereby offering more interpretable and standardized evaluations. These approaches demonstrate superior correlation with human judgments compared to other model-based metrics like COMET and BLEURT. Given that LLMs typically support longer context lengths, they are well-suited for evaluating tasks involving extended input sequences. Inspired by GEMBA-MQM (Kocmi and Federmann, 2023a), we employ GPT-4 as an evaluator to investigate the qualitative differences between baseline models.
@@ 3.1 Model
Our experimentation focuses on standard Encoder-Decoder transformer models. While several massively multilingual models, such as NLLB (Costa-jussà et al., 2022) and its distilled variants, have been released, our experimentation requires the availability of long-context or document-level training and evaluation datasets. Consequently, we focus on language-family specific models that allow for more controlled experimentation. For our experiments, we chose the distilled variants of IndicTrans2 Gala et al. (2023); Gumma et al. (2023). A key difference is that the distilled models released in Gala et al. (2023) we used were not trained with BackTranslation Sennrich et al. (2016) data, which we incorporate in our experimental setup. We experimented with models based on RoPE and ALiBi by replacing the Sinusoidal Positional Embedding in the pre-trained model with the corresponding relative Positional Embedding. This replacement led to a significant performance drop, prompting us to fine-tune the models using high-quality data Mohiuddin et al. (2022); Gumma et al. (2023); Gala et al. (2023) to study their recovery in each setting.
@@ 3.2.1 Train set
For maximum performance gains for our models, we choose a high-quality mixture of seen and unseen distributions for our training data. To this end, for the seen distribution, we select the BPCC-Seed, which is a high-quality human-verified subset of BPCC (Bharat Parallel Corpus Collection), and has shown to be effective in improving model performance Gala et al. (2023). As for the unseen distribution, we sample the top 150K pairs for each language from the BackTranslation (BPCC-BT) subset of BPCC based on the provided LASER Costa-jussà et al. (2022) or LaBSE Feng et al. (2022) cosine-similarity scores333whichever was available for that language. In case similarity scores are unavailable for a translation pair, a random sample was chosen. Additionally, for the unseen data, we also include the Asian Language Treebank (ALT) Corpus Riza et al. (2016), and the CoPara Aligned corpus E et al. (2023), which contains document-level translation pairs for four Dravidian languages, Telugu, Kannada, Malayalam and Tamil. The overall statistics of the dataset is available in table 1.
@@
Table 2: Distribution of long-context data in the fine-tuning data mixture. Indic-En pairs also follow the same distribution. The language-pairs not mentioned do not have any data in the aforementioned lengths.
@@ 3.2.2 Dev/Test sets
Most existing benchmarks for Indian languages are composed of sentence-level test sets, which may not adequately reflect long-context performance or extrapolation capabilities. To address this, we construct document-level development and test sets444 https://aka.ms/DocMTBenchmarks using Flores (Costa-jussà et al., 2022) and IN22-Conv (Gala et al., 2023). In this setup, sentences are grouped within the same context window, making them suitable for document-level evaluation. For Flores, we group sentences by metadata, including URL, domain, and topic headers, to map sentence-level pairs into document-level pairs, where each document-level unit consists of three sentences. This method is viable since this benchmark was created by professional translators who translate three consecutive sentences from every selected context window, as described by Goyal et al. (2022). As a result, we obtain 286 merged sentence pairs per language, with a total of 5720 pairs. In case of IN22-Conv, the dataset was originally created by translating entire conversations from English into the target languages by professional translators, and used for sentence-level evaluation by Gala et al. (2023). We use metadata provided to trace back the conversations and treat the entire conversation as a sequence for long-context evaluation. This dataset has 44 conversations per language and covers 22 languages. The token length distribution of the synthesized IN22-Conv test set is shown in Figure 1. IN22-Conv serves as an accurate representation of a true document-level evaluation set due to its contextual dependencies across multiple turns in the conversation. Additionally, it supports all the languages under consideration, making it the primary benchmark for document-level evaluation.
@@ 3.3 Metrics and Evaluation
For consistency, we use the ChrF++ score (Popović, 2017) as the primary string-based metric, and follow the same evaluation practices as outlined in Gala et al. (2023) by normalizing the indic texts using IndicNLPLibrary [^5] https://github.com/anoopkunchukuttan/indic_nlp_library. For brevity, we report the aggregate score across all the languages in the respective test set, which are denoted in Table 3. For GPT-4o-based evaluation, we use the prompt template from GEMBA-MQM (Kocmi and Federmann, 2023a) as a starting point and make slight adjustments to elicit an implicit quality score between 0 and 1, in addition to identifying the error spans following the MQM guidelines. Another notable difference is that, since we are evaluating long-context translations, including few-shot examples would be inefficient due to the limited context length. Therefore, we opt for a zero-shot evaluation setup. We report both the implicit score as rated by GPT-4o as well as calculated MQM-score (based on the scoring guidelines outlined in Kocmi and Federmann (2023a). We start with an ideal score of 25, deducting points based on identified errors and their severity: Critical errors are penalized by a factor of 5, Major by 3, and Minor by 1. The final score is normalized by dividing by 25, yielding a value between 0 and 1.
@@ 3.4 Finetuning Strategies
First, we choose a full-fine-tuning (FFT) approach which tweaks all the 211.78M parameters of the model. This would allow the embeddings, layer-normalizations, and the rest of the linear layers [^6] SelfAttention, CrossAttention, FeedForward to be tuned as per the new relative Positional Embedding.
@@ 4 Experiments
We use the fairseq [^7] Our variant implements, LoRA, RoPE, ALiBi, and a slightly efficient MultiHeadAttention as per the requirements of this work framework Ott et al. (2019) for all of our experiments. All the finetuning experiments were conducted on 8 $\times$ H100 80Gb GPUs, and every model was trained till convergence with early-stopping. The exact hyperparameter sets are mentioned in Table 4.
Notably, for RoPE, we apply the rotation function to only the first half segment of the query and key feature vectors while leaving the rest as is. We empirically notice that the partial-rotation [^8] https://github.com/lucidrains/x-transformers/issues/40 implementation gives slightly better performance as compared to a full-rotation.
@@
Patience 20
Max Positions 4096
Warmup steps 4000
@@ 5 Results and Discussions
Figure 2: ChrF++ scores on Sentence-Level (top) and Document-Level (bottom) benchmarks. Results are presented for baselines using three fine-tuning setups (FFT, LoRA, and min-LoRA) that compare three types of positional embeddings (Sine, ALiBi, and RoPE), along with the pre-trained model performance baseline.
@@ 5.1 Swapping out Positional Embeddings
Figure 2 demonstrate the performance of several experimental setups that involve swapping of Positional Embeddings and followed by fine-tuning (if applicable). The raw scores are presented in Table 6.
@@
A notable finding in this setting is that fine-tuning a model originally using Sinusoidal embeddings yielded performance comparable to models fine-tuned after the switch to using RoPE or ALiBi embeddings on sentence-level benchmarks as observed in Figure 2. This further reinforces the conclusion that Positional Embeddings in pre-trained transformer models can be safely tweaked even in later stages. Although Sinusoidal, RoPE, and ALiBi embeddings show similar performance in sentence-level test sets. The advantage of switching to RoPE or ALiBi can be observed in case of document-level test sets in line with the findings of Su et al. (2024) and Press et al. (2022).
@@
Our fine-tuning dataset primarily comprised sentence-level data, as detailed in Section 3.2.1, with only a limited amount of long-context or paragraph-level data available for four Dravidian languages accounting for less than 0.03% of the complete fine-tuning data mixture (see Table 1). The token distribution across this dataset is illustrated in Table 2. While most languages lacked long-context data, evaluation on document-level benchmarks still demonstrated strong performance on these, as seen in Figure 2. This indicates that fine-tuning with a small amount of long-context data introducing relative positional embeddings can enable robust cross-lingual length generalization. This finding is particularly important because long-context data is often scarce across many languages. Nevertheless, effective adaptation to longer contexts was achieved through cross-lingual generalization, even with minimal long-context data in select languages.
@@
RoPE has lower throughput compared to Sinusoidal embeddings due to its higher number of operations. Since the rotation function in RoPE is applied at every layer, it has a time complexity of $\mathcal{O}(n)$, whereas Sinusoidal embeddings are added only once, after the token embeddings, resulting in $\mathcal{O}(1)$ complexity. The exact decoding rates (in tokens per second) are illustrated in Figure 4. An efficient implementation of ALiBi Press et al. (2022) achieves $\mathcal{O}(1)$ complexity, making it slightly slower than Sinusoidal but still significantly faster than RoPE. Therefore, for a basic sentence-level MT system, Sinusoidal PEs offer the best efficiency in terms of throughput and performance. However, for document-level or longer contexts, RoPE provides notable qualitative advantages, although it comes with a decrease in throughput.